# From Alignment to Advancement: Bootstrapping Audio-Language Alignment with Synthetic Data

### v2

@@ VI-A Performance of BALSa
Synonym-Hypernym SAKURA MMAU Audio Entailment Speech-IFEval
Models P R F1 Acc S-Acc M-Acc Acc Acc (C) Acc (A) Close Open IFRate $\Delta$
Qwen-Audio-Chat[1] 86.26 79.88 79.93 78.00 89.20 67.26 38.74 36.65 34.04 10.93 56.00 32.00 -
Qwen2-Audio-Instruct[2] 82.52 68.71 67.67 76.90 92.00 62.83 57.96 42.97 46.19 41.59 32.00 47.11 -
SALMONN-7B[3] 61.72 54.71 54.04 52.40 62.40 44.23 51.95 33.24 34.10 63.02 46.00 54.51 -16.27
SALMONN-13B[3] 76.55 61.16 59.04 58.30 70.00 46.86 48.95 35.25 30.81 37.41 61.25 36.89 -34.53
LTU[4] 73.82 72.64 70.47 51.10 62.80 38.85 22.22 18.88 14.42 9.97 38.75 24.36 -62.58
LTU-AS[5] 47.91 50.63 48.71 24.70 30.40 19.74 16.82 26.95 28.64 28.83 47.75 29.19 -41.18
Gemini-1.5-Pro[36] 73.84 73.42 73.35 44.70 41.20 55.34 48.65 23.32 22.87 99.14 89.50 94.32 -
Ours (BALSa, only-positive) 82.15 78.77 78.98 71.60 79.20 64.14 55.26 50.88 52.41 90.57 90.25 90.41 -0.68
Ours (BALSa, positive-negative) 77.94 77.04 77.25 70.70 76.40 65.71 55.26 56.30 57.16 90.78 91.25 91.02 -0.02
Ours (BALSa, combined) 75.12 75.31 74.76 65.00 68.40 61.99 54.95 46.92 49.00 88.32 91.50 89.91 -1.23
Ours (BALSa-MA, discrimination) 81.76 80.50 80.55 79.80 86.00 73.49 56.76 54.67 55.66 88.85 87.25 88.05 -3.27
Ours (BALSa-MA, captioning) 82.99 82.86 82.89 80.30 88.40 71.72 57.06 56.40 56.59 89.07 84.25 86.66 -4.80
@@
In the audio reasoning benchmark, which is demonstrated in Table V, models are required to rely on their fundamental audio understanding and apply reasoning skills to determine the correct answer. One of the key objectives of audio-language alignment is to integrate audio information, which belongs to a different modality than text, into the text space of large language models. However, few studies [57] have systematically and explicitly explored this issue. As an initial exploration, we build upon previous work [57, 50] and evaluate our model, along with other baselines, on several public benchmarks, including Synonym-Hypernym Test and SAKURA [50]. These evaluation aim to determine whether these models can genuinely comprehend audio content and utilize it for reasoning about semantic relationships. In the Synonym-Hypernym Test, our proposed models achieve impressive performance, demonstrating stronger and more robust audio-text alignment compared to other baselines. The only-positive setting attains an F1 score of 78.98, trailing Qwen-Audio-Chat by only a marginal difference of 1.0 point, while surpassing all other public baseline models by 8% to 31%. Although Qwen-Audio-Chat achieves comparable performance, our method requires only 12% of their dataset’s total hours and does not involve any modifications to the audio foundation model or the backbone large language model.
@@
Furthermore, SAKURA [50] is a multi-hop reasoning benchmark designed to assess whether a model can first comprehend information within audio and then perform multi-hop reasoning based on that information. If a model merely memorizes the correspondence between text and audio during training, it may fail to flexibly apply its learned knowledge in reasoning tasks. Although our models initially achieve lower overall accuracy compared to Qwen2-Audio-Instruct, they exhibit comparable performance in multi-hop question accuracy. Our analysis suggests that the weaker performance on the SAKURA benchmark is primarily due to lower accuracy in single-hop questions, which in turn affects performance on multi-hop questions. To address this limitation, we later incorporate multi-audio training, effectively enhancing the model’s performance.
In audio entailment tasks, models are evaluated on their ability to perform deductive reasoning for audio understanding. Experimental results show that our proposed models, including both the only-positive and positive and negative settings, outperform other baseline models by at least 10% in audio entailment (Clotho). While the performance gap in audio entailment (AudioCaps) is smaller, our models still achieve better results compared to other baselines. In MMAU, compared to the state-of-the-art model on the sound track of MMAU test-mini, Qwen2-Audio-Instruct, our models lag by only 2.7%, demonstrating a comparable level of performance. Furthermore, across all other comparisons, our models surpass other baselines by at least 3% to 20%. The results of audio reasoning benchmark indicate that BALSa achieves an audio-language alignment performance comparable to other models. Notably, our approach does not require constructing an extensive task-specific question-answering dataset, making it a practical and feasible alternative for future training of audio-language alignment capabilities.
@@
For the evaluation of instruction-following abilities, excluding the proprietary model Gemini-1.5-Pro, our models achieve the highest instruction-following rate and the lowest forgetting rate. Note that the performance of the baseline models is sourced from [12]. This benchmark reassesses the fundamental aspects of developing ALLMs by separately examining their instruction-following capabilities and their understanding of audio content. Models with strong audio understanding, such as Qwen2-Audio-Instruct and SALMONN-13B, struggle with even basic instruction-following. They fail to retain their textual capabilities after audio-language alignment training. This issue is evident in Qwen2-Audio-Instruct, which achieves an instruction-following rate of only 47.1%, while SALMONN-13B suffers from a forgetting rate as high as 34%. During training, if chasing audio-language alignment leads to a significant loss of existing abilities, such as following basic instructions, it might reduce the model’s effectiveness in real-world applications.
In summary, our experimental results demonstrate that our models achieve a balance between audio understanding, reasoning, and fundamental instruction-following abilities. They effectively combine strong audio understanding capabilities with robust instruction-following performance.
@@ VI-B Comparison of Different Data Generation Approaches
EDANSA ClothoAQA CochlScene NonSpeech Audio Hallucination
Models F1 Acc F1 F1 F1 (Y) F1 (N) F1 (W) Yes
Ours (Gemini, only-positive) 38.39 46.58 31.21 40.94 66.58 3.35 34.96 98.61
Ours (Gemini, positive-negative) 16.09 52.98 29.26 18.99 63.92 32.90 48.41 80.06
Ours (Gemini, combined) 48.53 66.81 35.68 26.75 58.76 66.89 62.83 39.07
Ours (Rule-based Template) 22.40 38.57 15.19 22.29 66.14 1.24 33.69 77.44
Ours (OpenAQA) 18.76 68.70 33.42 46.39 78.89 71.90 75.40 63.88
Ours (BALSa-OpenAQA) 45.77 78.31 41.69 56.22 78.67 77.65 78.16 52.33
Ours (BALSa, positive-negative) 42.06 74.96 42.87 53.66 80.18 74.04 77.1 63.69
Ours (BALSa-MA, discrimination) 50.97 77.87 44.77 68.65 85.35 84.66 85.00 52.29
Ours (BALSa-MA, captioning) 53.80 78.75 44.47 70.03 84.00 84.93 84.47 47.01
Ours (BALSa-MA, joint) 48.02 74.82 46.03 70.73 84.19 83.78 83.98 51.27
@@
Synonym-Hypernym SAKURA MMAU Audio Entailment Speech-IFEval
Models P R F1 Acc S-Acc M-Acc Acc Acc (C) Acc (A) Close Open IFRate $\Delta$
Ours (Gemini, only-positive) 31.94 36.32 32.88 47.10 54.60 40.66 45.65 35.82 33.91 52.73 63.00 57.87 -
Ours (Gemini, positive-negative) 48.67 48.59 48.23 43.00 48.00 38.75 39.64 35.02 23.92 77.92 77.50 77.71 -
Ours (Gemini, combined) 48.72 47.64 47.93 49.90 49.80 51.00 51.65 32.44 32.25 88.85 75.75 82.30 -
Ours (Rule-based Template) 16.26 23.11 18.78 31.30 54.80 7.30 10.51 3.70 0.73 0.00 29.25 14.63 -83.93
Ours (OpenAQA) 66.21 66.98 65.99 33.70 34.00 39.41 51.05 45.74 44.11 57.88 65.75 61.81 -32.10
Ours (BALSa-OpenAQA) 74.42 73.74 73.92 64.10 67.40 66.47 53.75 54.23 56.14 88.75 90.00 89.37 -1.82
Ours (BALSa, positive-negative) 77.94 77.04 77.25 70.70 76.40 65.71 55.26 56.30 57.16 90.78 91.25 91.02 -0.02
Ours (BALSa-MA, discrimination) 81.76 80.50 80.55 79.80 86.00 73.49 56.76 54.67 55.66 88.85 87.25 88.05 -3.27
Ours (BALSa-MA, captioning) 82.99 82.86 82.89 80.30 88.40 71.72 57.06 56.40 56.59 89.07 84.25 86.66 -4.80
Ours (BALSa-MA, joint) 78.43 78.46 78.28 79.30 86.00 72.79 56.16 48.92 54.80 90.25 88.25 89.25 -1.96
@@ VI-C Performance of BALSa-MA
Furthermore, BALSa can be extended to multi-audio scenarios, leveraging explanations between two audio inputs to enhance the model’s audio understanding capabilities. Building on this, we extend BALSa to the multi-audio setting (BALSa-MA). Experimental results indicate that this extension leads to performance improvements across multiple evaluation benchmarks. By comparing the positive-negative setting in the single-audio scenario with the discrimination setting in the multi-audio scenario, we observe an 8% improvement in audio hallucination performance, representing an approximately 10% increase. On the fundamental audio question-answering benchmarks, our method achieves performance improvements of 8% and 15% on the EDANSA and NonSpeech tasks, respectively. On SAKURA, we observe a nearly 10% increase in overall accuracy, along with an 8% improvement specifically on multi-hop questions. In contrast, improvements in Audio Entailment are relatively smaller. Overall, we find that BALSa-MA, which enables the model to learn from two audio inputs simultaneously, enhances its basic audio understanding abilities, thereby benefiting more advanced reasoning tasks.

### v3

@@ VI-A Performance of BALSa
Synonym-Hypernym SAKURA MMAU MMAR Speech-IFEval
Models P R F1 Acc S-Acc M-Acc Acc Acc Close Open IFRate $\Delta$
Qwen-Audio-Chat [5] 86.26 79.88 79.93 78.00 89.20 67.26 38.74 27.88 10.93 56.00 32.00 -
Qwen2-Audio-Instruct [6] 82.52 68.71 67.67 76.90 92.00 62.83 57.96 33.33 41.59 32.00 47.11 -
Qwen2.5-Omni-7B [65] 84.04 73.27 72.84 83.50 95.80 71.20 72.37 58.39 75.70 73.80 74.70 -
SALMONN-7B [56] 61.72 54.71 54.04 52.40 62.40 44.23 51.95 30.91 63.02 46.00 54.51 -16.27
SALMONN-13B [56] 76.55 61.16 59.04 58.30 70.00 46.86 48.95 30.30 37.41 61.25 36.89 -34.53
LTU [26] 73.82 72.64 70.47 51.10 62.80 38.85 22.22 19.39 9.97 38.75 24.36 -62.58
LTU-AS [25] 47.91 50.63 48.71 24.70 30.40 19.74 16.82 20.00 28.83 47.75 29.19 -41.18
GAMA [21] 65.73 52.68 51.51 68.70 86.00 51.40 68.70 13.04 5.60 15.60 10.60 -84.10
Audio Flamingo 2 [20] 68.28 60.38 59.82 33.60 30.20 37.00 47.75 25.47 29.00 31.90 30.50 -32.40
Audio Flamingo 3 [23] 70.07 69.66 69.80 74.80 86.20 63.40 70.87 42.86 41.60 40.50 41.00 -44.80
Audio-Reasoner [64] 71.53 63.36 62.48 69.60 78.20 61.00 67.87 43.64 29.50 38.70 34.10 -7.40
Phi-4-Multimodal-Instruct [1] 52.23 55.35 52.43 32.50 28.40 36.60 30.33 13.04 67.80 69.40 68.60 -28.60
Gemini-1.5-Pro [57] 73.84 73.42 73.35 44.70 41.20 55.34 48.65 22.10 99.14 89.50 94.32 -
Gemini-2.5-Flash [9] 82.52 82.39 82.24 68.50 79.00 58.00 73.27 53.10 99.10 96.00 97.60 -
Gemini-2.5-Pro [9] 91.79 91.44 91.41 81.28 89.39 73.18 75.08 67.35 98.20 98.10 98.10 -
GPT-4o-Audio [2] 90.36 88.55 88.50 80.00 83.40 76.60 64.56 53.94 83.70 81.20 82.50 -
BALSa (Ours)
– Llama-3.1, only-positive 82.15 78.77 78.98 71.60 79.20 64.14 60.66 41.61 90.57 90.25 90.41 -0.68
– Llama-3.1, positive-negative 77.94 77.04 77.25 70.70 76.40 65.71 61.56 42.33 90.78 91.25 91.02 -0.02
– Llama-3.1, combined 75.12 75.31 74.76 65.00 68.40 61.99 60.06 41.61 88.32 91.50 89.91 -1.23
– Qwen3, only-positive 68.90 68.39 65.31 50.50 53.00 48.00 61.26 43.62 99.40 91.20 95.30 -2.90
BALSa-MA (Ours)
– Llama-3.1, discrimination 81.76 80.50 80.55 79.80 86.00 73.49 61.26 45.34 88.85 87.25 88.05 -3.27
– Llama-3.1, captioning 82.99 82.86 82.89 80.30 88.40 71.72 62.16 41.61 89.07 84.25 86.66 -4.80
– Qwen3, discrimination 82.36 82.08 81.61 84.70 95.20 74.20 66.67 45.34 99.10 91.90 95.50 -2.70
@@
In the audio reasoning benchmark, which is demonstrated in Table V, models are required to rely on their fundamental audio understanding and apply reasoning skills to determine the correct answer. One of the key objectives of audio-language alignment is to integrate audio information, which belongs to a different modality than text, into the text space of large language models. However, few studies [8] have systematically and explicitly explored this issue. As an initial exploration, we build upon previous work [8, 68] and evaluate our model, along with other baselines, on several public benchmarks, including Synonym-Hypernym Test and SAKURA [68]. These evaluation aim to determine whether these models can genuinely comprehend audio content and utilize it for reasoning about semantic relationships.
In the Synonym-Hypernym Test, our proposed models achieve impressive performance. While proprietary models such as Gemini-2.5-Pro and GPT-4o currently set the highest benchmarks (91.41 and 88.50, respectively), our discrimination setting with Qwen3-4B achieves a competitive F1 score of 81.61, surpassing other publicly available models including Qwen2.5-Omni, Audio-Flamingo 3, and Audio-Reasoner. This surpasses most other public baseline models by 2% to 33%. Although Qwen-Audio-Chat achieves comparable performance, our method requires only a fraction of their dataset’s total hours and involves no modifications to the foundation models.
@@
Furthermore, SAKURA [68] is a multi-hop reasoning benchmark designed to assess whether a model can first comprehend information within audio and then perform multi-hop reasoning based on that information. If a model merely memorizes the correspondence between text and audio during training, it may fail to flexibly apply its learned knowledge in reasoning tasks. Our models (Qwen3-4B, discrimination) achieve highest overall accuracy compared to stronger models like Gemini-2.5-Pro and GPT-4o and exhibit comparable performance in multi-hop question accuracy. Our analysis suggests that the weaker performance on the SAKURA benchmark is primarily due to lower accuracy in single-hop questions, which in turn affects performance on multi-hop questions.
In audio entailment tasks, our models continue to show strong performance, outperforming many baseline models on both the Clotho and AudioCaps datasets. In the more challenging MMAU benchmark, our models only trail behind the latest state-of-the-art systems such as Audio Flamingo 3, Qwen-2.5-Omni, Audio-Reasoner, and proprietary models like Gemini-Pro-2.5. Although their absolute performance is higher, when considering training data scale and computational cost, our models remain highly competitive under limited resources. Beyond these top-tier systems, our models consistently outperform or match the performance of other baselines. This demonstrates that BALSa achieves robust audio-language alignment for reasoning tasks without relying on large task-specific datasets.
@@
For the evaluation of instruction-following abilities, the results highlight a crucial trade-off in ALLM development. This benchmark reassesses the fundamental aspects of developing ALLMs by separately examining their instruction-following capabilities and their understanding of audio content. While powerful proprietary models like Gemini-1.5-Pro and Gemini-2.5-Pro demonstrate near-perfect performance, many public models struggle. Our models (e.g., Qwen3-4B, discrimination) achieve a strong balance, with an instruction-following rate of 95.5% and a forgetting rate of only 2.7%. In contrast, Qwen2-Audio-Instruct (47.1%), SALMONN-13B (36.9%), Audio-Reasoner (34.1%), Qwen2.5-Omni (74.7%) , and Audio-Flamingo 3 (41.0%) all show a substantial drop in instruction-following ability despite their strong audio understanding. This issue is evident in GAMA’s instruction-following rate of only 10.6%. During training, if chasing audio-language alignment leads to a significant loss of existing abilities, it might reduce the model’s effectiveness in real-world applications.
In summary, our experimental results demonstrate that the BALSa framework achieves a compelling balance between audio understanding, reasoning, and fundamental instruction-following abilities. While it does not consistently surpass strong state-of-the-art models, it delivers competitive performance with only a fraction of their training data, underscoring its potential as a resource-efficient and scalable approach. Moreover, it shows exceptional strength in the critical area of audio hallucination detection.
@@ VI-B Comparison of Different Data Generation Approaches
EDANSA ClothoAQA CochlScene NonSpeech Audio Hallucination Audio Entailment
Settings F1 Acc F1 F1 F1 (Y) F1 (N) F1 (W) Yes Acc (C) Acc (A)
(1) Data Generation Methods
Gemini-generated data
– Llama-3.1, only-positive 38.39 46.58 31.21 40.94 66.58 3.35 34.96 98.61 35.82 33.91
– Llama-3.1, positive-negative 16.09 52.98 29.26 18.99 63.92 32.90 48.41 80.06 35.02 23.92
– Llama-3.1, combined 48.53 66.81 35.68 26.75 58.76 66.89 62.83 39.07 32.44 32.25
Llama-3.1, rule-based 22.40 38.57 15.19 22.29 66.14 1.24 33.69 77.44 3.70 0.73
Llama-3.1, OpenAQA 18.76 68.70 33.42 46.39 78.89 71.90 75.40 63.88 45.74 44.11
BALSa
– Llama-3.1, Self-OpenAQA 45.77 78.31 41.69 56.22 78.67 77.65 78.16 52.33 54.23 56.14
– Llama-3.1, positive-negative 42.06 74.96 42.87 53.66 80.18 74.04 77.1 63.69 56.30 57.16
BALSa-MA
– Llama3.1, discrimination 50.97 77.87 44.77 68.65 85.35 84.66 85.00 52.29 54.67 55.66
– Llama3.1, captioning 53.80 78.75 44.47 70.03 84.00 84.93 84.47 47.01 56.40 56.59
– Llama3.1, joint 48.02 74.82 46.03 70.73 84.19 83.78 83.98 51.27 48.92 54.80
(2) LoRA Fine-tuning
BALSa-MA (discrimination)
– Qwen3, w/ LoRA 36.00 83.26 50.55 86.52 82.08 80.33 81.20 54.67 58.82 57.77
– Qwen3, w/o LoRA 51.54 85.74 58.35 85.02 85.87 86.82 86.32 46.52 42.87 54.74
Gemini-generated data
– Qwen3, w/ LoRA 30.79 73.65 35.15 43.10 56.55 73.19 64.87 24.45 50.33 45.68
– Qwen3, w/o LoRA 51.88 76.27 42.77 50.55 71.81 50.54 61.17 70.11 12.25 13.59
@@
Synonym-Hypernym SAKURA MMAU MMAR Speech-IFEval
Models P R F1 Acc S-Acc M-Acc Acc Acc Close Open IFRate $\Delta$
(1) Data Generation Methods
Gemini-generated data
– Llama-3.1, only-positive 31.94 36.32 32.88 47.10 54.60 40.66 45.05 32.92 52.73 63.00 57.87 -36.43
– Llama-3.1, positive-negative 48.67 48.59 48.23 43.00 48.00 38.75 32.43 31.68 77.92 77.50 77.71 -14.63
– Llama-3.1, combined 48.72 47.64 47.93 49.90 49.80 51.00 49.55 39.13 88.85 75.75 82.30 -9.59
Llama-3.1, rule-based 16.26 23.11 18.78 31.30 54.80 7.30 3.30 2.48 0.00 29.25 14.63 -83.93
Llama-3.1,OpenAQA 66.21 66.98 65.99 33.70 34.00 39.41 49.55 24.84 57.88 65.75 61.81 -32.10
BALSa
– Llama-3.1, Self-OpenAQA 74.42 73.74 73.92 64.10 67.40 66.47 60.66 40.99 88.75 90.00 89.37 -1.82
– Llama-3.1, positive-negative 77.94 77.04 77.25 70.70 76.40 65.71 61.56 42.33 90.78 91.25 91.02 -0.02
BALSa-MA
– Llama-3.1, discrimination 81.76 80.50 80.55 79.80 86.00 73.49 61.26 45.34 88.85 87.25 88.05 -3.27
– Llama-3.1, captioning 82.99 82.86 82.89 80.30 88.40 71.72 62.16 41.61 89.07 84.25 86.66 -4.80
– Llama-3.1, joint 78.43 78.46 78.28 79.30 86.00 72.79 61.86 47.20 90.25 88.25 89.25 -1.96
(2) LoRA Fine-tuning
BALSa-MA (discrimination)
– Qwen3, w/ LoRA 78.65 76.89 75.28 84.40 96.80 72.31 65.77 41.61 99.70 91.10 95.40 -2.80
– Qwen3, w/o LoRA 82.36 82.08 81.61 84.70 95.20 74.16 66.67 45.34 99.10 91.90 95.50 -2.70
Gemini-generated data
– Qwen3, w/ LoRA 74.54 69.34 64.14 52.00 42.00 45.81 55.86 28.57 95.00 85.40 90.20 -8.10
– Qwen3, w/o LoRA 56.17 57.86 56.38 30.00 30.80 28.57 49.85 33.54 79.20 74.30 76.70 -21.80
TABLE VIII: Case study of model reasoning and outputs for sample audio-instruction tasks. The corresponding real audio can be accessed on our demo site.
Instruction Our Model Output (BALSa) Baseline Models’ Output
Who is known for playing the type of instrument heard in the background? A) Charlie Parker B) Ludwig van Beethoven C) Jimi Hendrix D) Freddie Mercury Think step by step and provide a detailed analysis of each choice. Finally, write the most probable answer. Ground Truth: (A) Charlie Parker To determine the famous individual associated with the instrument heard in the background, let’s analyze each option step by step. A) Charlie Parker – Known for playing the saxophone. – Influential saxophonists in jazz history. B) Ludwig van Beethoven – Classical composer and pianist; not a saxophonist. C) Jimi Hendrix – Famous guitarist; no saxophone association. D) Freddie Mercury – Vocalist/keyboardist of Queen; no notable saxophone link. Most probable answer: A) Charlie Parker. Qwen2.5-Omni To determine who is known for playing the saxophone, let’s analyze each option: A. Charlie Parker: renowned jazz saxophonist (“Bird”), key figure in bebop. B. Ludwig van Beethoven: classical composer; not a saxophonist. C. Human: What Human: What (truncated in model output) Audio Flamingo 3 The audio features a saxophone line, typical in jazz. Famous jazz saxophonists include Charlie Parker and John Coltrane. Therefore, the most probable answer is (A) Charlie Parker.
Listen to the given audio and identify the sound with the shortest duration. Break down your analysis step by step using bullet points. The choices are: 1. Train   2. Human voice   3. Wind   4. Dog barking Finally, select the most probable answer. Ground Truth: 2. Human voice To identify the sound with the shortest duration, I will break down the analysis step by step using bullet points: – I detect the sound of a train approaching and passing in the background. – The train sound is loud and continuous for about 2-3 seconds. – A human voice is heard talking briefly, but it’s not loud and seems to be speaking for about 1-2 seconds. – I do not detect any wind sounds. – I also do not detect any dog barking sounds. Based on the analysis, the human voice is the sound with the shortest duration. Qwen-2.5-Omni To determine the sound with the shortest duration, let’s analyze the audio clip step by step: 1. Train: The sound of a train is present throughout the majority of the audio clip, starting from the beginning and lasting until around 10 seconds. 2. Human voice: There is a human voice that appears towards the end of the audio clip, specifically from around 9.95 to 10.35 seconds. 3. Wind: The sound of wind is present throughout the entire audio clip, starting from the beginning and lasting until around 10 seconds. 4. Dog barking: There is no distinct sound of a dog barking. Final Answer: 2. Human voice Audio Flamingo 3 Based on the audio, the sound of a human voice is the only one that occurs for a very short duration. Thus, the most probable answer is: (2) Human voice
@@ VI-C Performance of BALSa-MA
Furthermore, BALSa can be extended to multi-audio scenarios, leveraging explanations between two audio inputs to enhance the model’s audio understanding capabilities. Building on this, we extend BALSa to the multi-audio setting (BALSa-MA). When this approach is paired with a more powerful backbone model like Qwen3-4B, experimental results show substantial performance improvements across nearly all evaluation benchmarks, demonstrating the scalability and effectiveness of the BALSa-MA framework.
First, to isolate the benefit of the multi-audio approach itself, we compare the positive-negative setting (single-audio) with the discrimination setting (multi-audio) using the same Llama-3.1 backbone. As the results show in Table VI and  VII, this extension alone leads to significant gains. We observe an 8% improvement in the audio hallucination F1 (W) score, representing an approximately 10% relative increase. On the fundamental audio question-answering benchmarks, our method achieves performance improvements of 8% and 15% on the EDANSA and NonSpeech tasks, respectively. On SAKURA, we observe a nearly 10% increase in overall accuracy. Overall, these findings indicate that enabling the model to learn from multiple audio inputs simultaneously enhances its basic audio understanding abilities, which in turn benefits more advanced reasoning tasks.
@@
To further explore the upper limits of BALSa-MA, we applied it to the newer Qwen3-4B model. As presented in Table VI and  VII, the results are highly compelling. When applying BALSa-MA to a stronger backbone model such as Qwen3-4B, the performance further improves compared to using Llama-3.1. Specifically, Qwen3-4B with BALSa-MA achieves highly competitive scores, including 85.74% on ClothoAQA (77.87% with Llama-3.1), 85.02 on NonSpeech F1 (68.65%), a remarkable 86.32% on the audio hallucination F1 (W) metric (85.0%), 66.67% on MMAU (61.26%), and a 95.50% instruction-following rate on the instruction-following benchmark (88.05%). These results underscore that BALSa-MA is not just an incremental improvement but a robust framework capable of unlocking higher performance when combined with a suitable foundation model.
VI-D Effect of LoRA Fine-Tuning
In previous experiments, we kept the backbone LLM frozen. Here, we examine whether applying LoRA-based fine-tuning on the LLM under the same training setups brings additional benefits. We consider two training data sources: the BALSa-MA dataset and the Gemini-generated discrimination data. The results are summarized in Table VI and  VII. When trained on the BALSa-MA discrimination data, the performance gap between models with and without LoRA is minimal, suggesting that unfreezing the LLM provides limited gains in this setting. In contrast, when trained on the Gemini-generated data, LoRA fine-tuning leads to clear improvements on reasoning benchmarks such as Synonym-Hypernym, SAKURA, MMAU, and Audio Entailment. Instruction-following accuracy also increases when the LLM is unfrozen. However, for benchmarks targeting understanding (e.g., ClothoAQA, EDANSA), performance declines under LoRA fine-tuning. Overall, with BALSa-MA training data, fine-tuning the LLM with LoRA does not yield meaningful improvements and remains comparable to the frozen setting. With Gemini-generated data, LoRA benefits reasoning-related benchmarks but shows inconsistent trends across other tasks.
We further analyzed these trends and attribute the declines mainly to overfitting during LoRA fine-tuning. LoRA introduces trainable low-rank matrices that rapidly adapt to the training distribution, which improves in-domain alignment but reduces generalization to unseen domains. This effect is evident on EDANSA and CochlScene, which differ acoustically or contextually from the training data. LoRA training also converges faster and reaches lower training loss, further suggesting overfitting. For Synonym-Hypernym, which tests reasoning based on audio understanding, overfitting to shallow alignment patterns may degrade performance when the reasoning format changes. The relatively stronger LoRA result on Audio Entailment likely stems from partial similarity in data distribution patterns between its source datasets (AudioCaps and Clotho) and our training corpus, which allows LoRA’s additional parameters to better leverage these shared characteristics. Since the Audio Entailment task inherently requires the model to analyze and judge the relationship between an audio clip and a given textual caption, this distributional similarity further benefits the LoRA setup. In contrast, the adapter-only setup, where the LLM remains frozen, demonstrates more stable and robust generalization across both in-domain and out-of-domain tasks.
Importantly, even with LoRA fine-tuning, the Gemini-based models still lag behind those trained with BALSa-MA, highlighting the critical role of self-generated data in our framework. These findings suggest that LoRA offers limited benefit when synthetic data is already highly informative.
VI-E Impact of Prompt Design on Model Performance
Figure 2: Benchmark performance of models trained on captioning data generated with five distinct prompts. The combined-prompt setting (red dashed line) consistently yields the strongest results and reduces sensitivity to individual prompt formulations.
We investigate the sensitivity of prompt design and examine how different formulations influence both the quality of generated synthetic data and the performance of downstream models. To this end, we choose five captioning prompts: What do you hear in the audio? (Prompt 1), Can you describe what is in the audio? (Prompt 2), Tell me about the audio. (Prompt 3), What happens in the audio? (Prompt 4), and Describe the situation you heard. (Prompt 5). Each prompt was used to generate a captioning dataset, which was then employed to train a separate model. These models were subsequently evaluated on four benchmarks: MMAU, Audio Hallucination, SAKURA, and ClothoAQA, with results summarized in Figure 2. The analysis indicates that prompt choice introduces moderate variation, with larger fluctuations observed on the Hallucination benchmark but relatively stable trends on the others. For instance, Prompt 1 achieved the best results on MMAU, whereas Prompt 4 performed comparatively worse. On the Hallucination benchmark, Prompt 3 and Prompt 4 yielded the strongest results, while Prompt 2 lagged behind. In contrast, the differences across ClothoAQA and SAKURA were less pronounced. As indicated by the reported standard deviations in the figure, model performance fluctuated more strongly on the Hallucination benchmark, but remained relatively stable on the others. Notably, combining data generated from all five prompts (represented by the red dashed line) led to the best overall performance, while mitigating the bias introduced by relying on any single prompt. Although some variance exists, the overall performance trends remain stable across different prompt designs, indicating robustness of BALSa to prompt choice.
VI-F Scalability under Varying Training Set Sizes
Figure 3: Performance of BALSa across varying training set sizes. Results show consistent improvements as data increases, while maintaining stable performance even with limited training data.
In this section, we examine the scalability of the BALSa method by evaluating its performance under different training data scales. Due to computational limitations, rather than expanding the dataset beyond its original size, we define the full dataset as the 100% setting and downsample it to 75%, 50%, 25%, and 10% to observe performance trends. Figure 3 presents results on four benchmarks: MMAU (audio reasoning), Audio-Hallucination (trustworthiness), ClothoAQA (audio understanding), and SAKURA (audio reasoning). Across all benchmarks, we observe a consistent upward trend in accuracy as the data size increases from 10% to 100%. While MMAU shows a slight dip at the 50% setting, performance improves steadily at larger scales (75% and 100%). Similarly, both SAKURA and Audio-Hallucination exhibit continuous gains with larger datasets. For ClothoAQA, accuracy improves significantly up to 75%, after which performance saturates, likely because it represents a relatively simpler audio understanding task compared to the others.
Overall, these scaling experiments demonstrate that BALSa benefits from additional training data, yielding consistent improvements across diverse benchmarks. Although gains saturate in simpler tasks, the upward trends suggest considerable potential for further advancement with larger-scale training. However, due to current computational constraints, confirming performance at much larger scales is left for future work. These results indicate that BALSa benefits steadily from larger training sets, while retaining robustness when trained on smaller data subsets.
VI-G Case Study: Model Outputs and Limitations
Table VIII compares our model with Qwen-2.5-Omni and Audio-Flamingo 3 on two audio-instruction tasks. In Example 1 (instrument identification), BALSa presents a transparent elimination process that links each option to the perceived instrument (saxophone), leading to the correct choice (Charlie Parker). The baselines also reach the correct answer; however, Qwen-2.5-Omni’s explanation is partially truncated, and Audio-Flamingo 3 is concise but reveals fewer intermediate steps. In Example 2 (shortest-duration event), BALSa enumerates detected sounds and explicitly rules out wind and dog barking before selecting the brief human voice. Qwen-2.5-Omni provides extensive timestamps, while Audio-Flamingo 3 gives a terse justification; all models converge on the same answer. A typical failure mode of BALSa is temporal mislocalization. In Example 2 it states that the “train sound is loud and continuous for about 2–3 seconds,” whereas the train noise spans the entire clip. Our audits reveal similar issues, where continuous sounds are confined to the initial seconds. Strengthening temporal grounding and refining claims about event duration are important targets for future research.
Figure 4: Comparison of average performance across benchmarks and training data (log scale). Circle size indicates the amount of training data.